# Controlling superradiant phase transition in quantum Rabi model

### v1

@@ Abstract
In the ultrastrong-coupling regime, the quantum Rabi model can exhibit quantum phase transition (QPT) when the ratio of the qubit transition frequency to the frequency of the cavity field approaches infinity. However, it is challenging to control the QPT in few-body systems because of the limited coupling strength and the $A^{2}$ terms. Here, we propose a practical scheme to manipulate the QPT of quantum Rabi model in the strong-coupling regime. By applying a periodic frequency modulation to the two-level system in a standard quantum Rabi model in the strong-coupling regime, an anisotropic quantum Rabi model with ultrastrong and tunable coupling strengths for rotating and counter-rotating terms is obtained. The ground-state and excitation energy of this model in terms of the modulation parameters are studied. We find that the QPT of quantum Rabi model can be observed in the strong-coupling regime and externally controlled by the modulation.
@@ I Introduction
Quantum phase transition (QPT) QPT1999 has long been an important and fundamental phenomenon in many-body systems, and has attracted remarkable amount of study interest due to its special features. The QPT occurs when the non-thermal parameter, such as coupling strength, exceeds a critical value at zero temperature QPT1999 . For example, the superradiant phase transition occurs in the Dicke model Dicke when the light-matter coupling strength exceeds a critical value in the thermodynamic limit, i.e., $N\rightarrow\infty$ 1973dickle ; 1973spt ; Sunchangpu ; liyong ; CEPRL ; NLPRL ; ABPRL ; PRL120183603 . The control of the QPT in Dicke model has been discussed ABPRL ; 2023Modulation . Recently, it is found that the QPT can also occur in few-body systems, such as the quantum Rabi model (QRM) 1936Rabi ; 1937Rabi when the ratio of the qubit transition frequency $\omega_{0}$ to the frequency $\omega_{c}$ of the cavity field tends to infinity, i.e., $\eta=\omega_{0}/\omega_{c}\rightarrow\infty$ Hwang2015Rabi . Hereafter, many efforts have been devoted to study the QPT in QRM both theoretically and experimentally trapped ions ; nuclear ; ZhengShiBiao , such as the universal scaling and critical exponents MaoXinLiu2017 , QPT in the quantum Rabi triangle Rabi triangle ; Rabi lattice , dissipative phase transition in the open QRM dispassive QRM ; dar ; sar , QPT in the anisotropic QRM yingzujian ; PRAaqrm and quench dynamics in the anisotropic QRM L. T. ShenPRA . Even the QPT in QRM in the fully quantum limit has also been predicted full quantum limit .
@@
Though the QPT in QRM has been experimentally observed in a single trapped ion trapped ions and a nuclear magnetic resonance quantum simulator nuclear , it is still challenging to manipulate the ground-state QPT due to the limited coupling strength with the classical field (dispersion) limit $\eta\rightarrow\infty$ and the $A^{2}$ terms in cavity quantum electrodynamics (QED) A2 term ; Hwang2015Rabi ; no-go theorm . The $A^{2}$ terms can prevent the occurrence of the QPT in QRM Hwang2015Rabi , which can not be neglected in the ultrastrong-coupling regime in cavity QED system.
@@
In this work, we propose a practical scheme to manipulate the QPT in QRM. Specifically, a periodic modulation was applied to the qubit transition frequency in a standard Rabi model 2017manipulation ; C. Liu in the strong-coupling regime. Then an anisotropic QRM is generated with rotating and counter-rotating coupling strengths allowed to be tuned. Through the modulation, both the effective rotating and counter-rotating (CR) coupling strengths can be varied in a wide range and reach ultrastrong-coupling regime. The ground state is allowed to be in superradiant phase. To investigate the QPT of the anisotropic QRM, we calculated excitation energy and the ground-state phases in both the normal and superradiant phase in the whole coupling parameter space. We find that both the normal and superradiant phase can be reached in the strong-coupling regime via the modulation. We also find that Goldstone modes Hwang2016JC ; Goldstone exist in the superradiant phase. The dependence of excitation energy and ground-state phases on the modulation parameters are analyzed. The feature of the QPT is also investigated by calculating the derivatives of the ground-state energy. Our results show that the QPT in the QRM can be observed and manipulated externally in the strong-coupling regime. These results may inspire other studies on QPT engineering in other few-body systems.
@@
This paper is organized as follows. In Sec. II, we introduce a periodic modulation to the qubit transition frequency in a standard QRM and obtain an effective anisotropic QRM whose rotationg and counter-rotating terms are tunable. In Sec. III, we analyze the dependence of the effective QPT boundary and coupling strengths on modulation parameters. We obtain the excitation energies for five phases and the phase boundary in Sec. IV. In Sec. V, we figure out the phase diagram and verify the phase diagram by analyzing the order parameters. The feature of QPT is also discussed. We also discuss how to manipulate the ground-states phases by varying the modulation parameters in Sec. VI. In Sec. VII, we discuss the $A^{2}$ terms undsr modulation. Finally, we present a brief conclusion and discussions in Sec. VIII.
@@ II Model and Hamiltonian
The system we consider is composed by a two-level system (qubit) coupled to a cavity field, as shown in Fig. 1, which is described by the quantum Rabi model 1936Rabi ; 1937Rabi . The corresponding Hamiltonian reads
@@
is the usual Jaynes-Cummings (JC) Hamiltonian, and
@@
describes the CR interactions. Here, $\sigma_{\pm}=(\sigma_{x}\pm i\sigma_{y})$/2 with $\sigma_{x,y,z}$ being the Pauli matrices, and $a\ (a^{\dagger})$ is the annihilation (creation) operator of the cavity field with frequency $\omega_{c}$. We use $\left|e\right\rangle$ and $\left|g\right\rangle$ to label the excited and ground states of the qubit, respectively. The transition frequency of the qubit is $\omega_{0}$ and the coupling strength between the qubit and the cavity field is $g$.
@@
where $\xi$ and $\nu$ are the dimensionless modulation amplitude and frequency, respectively. After including the two-photon $A^{2}$ terms
@@
[…]
where $\mathcal{T}$ denotes the time-ordering operator, the Hamiltonian becomes
@@
where we have omitted the constant term in $H_{A^{2}}$. Here, $\delta_{n}=\omega_{0}-\omega_{c}^{\prime}+n\nu$ is the oscillating frequency for the $n$th sideband in the rotating terms, and $\Delta_{m}=\omega_{0}+\omega_{c}^{\prime}+m\nu$ is the oscillating frequency for the $m$th sideband in the CR terms. In the derivation of $\tilde{H}(t)$, we have used the Jacobi-Anger expansion
$e^{i\xi\sin(\nu t)}=\sum_{n=-\infty}^{\infty}J_{n}(\xi)e^{in\nu t}.$ (10)
Here, $J_{n}(\xi)$ is the $n$th Bessel function of the first kind. In Eq. (9), a rotating sideband $n_{0}$ and a CR sideband $m_{0}$ satisfying $\left|gJ_{n_{0}}(\xi)/\delta_{n_{0}}\right|=\max\{\left|gJ_{n}(\xi)/\delta_{n}\right|,n\in Z\}$ and $\left|gJ_{m_{0}}(\xi)/\Delta_{m_{0}}\right|=\max\{\left|gJ_{m}(\xi)/\Delta_{m}\right|,m\in Z\}$ can be selected. Here, “Z” denotes all integers. The oscillating frequencies of all rotating and CR sidebands can be written as $\delta_{n_{0}+s}=\delta_{n_{0}}+s\nu$ and $\Delta_{m_{0}+s}=\Delta_{m_{0}}+s\nu$ with $s$ an integer, respectively. Thus the other sidebands are separated from $n_{0}$ ($m_{0}$) sideband by $s\nu$. Under the conditions
@@
all the sidebands other than the $n_{0}$($m_{0}$) sideband can be discarded by using RWA. With $g_{A^{2}}\ll g\ll\omega_{0}$ and $g_{A^{2}}\ll 2\omega_{c}^{\prime}$, the $a^{2}$ and $a^{\dagger 2}$ terms with oscillating frequencies ±$\pm 2\omega_{c}^{\prime}$ from the two-photon $A^{2}$ Hamiltonian can also be neglected using RWA. To be focus, the validity of this approximation is discussed in detail in Sec. VII. The Hamiltonian $\tilde{H}(t)$ then can be approximated as
@@
where we have introduced the effective coupling strengths
$g_{r}(\xi)=gJ_{n_{0}}(\xi),\ g_{cr}(\xi)=gJ_{m_{0}}(\xi).$ (13)
The Hamiltonian describes an effective standard Rabi model with $g_{r}=g_{cr}$ or an effective anisotropic QRM with $g_{r}\neq g_{cr}$ QTxie in interaction picture. Note that, we can tune the effective rotating and CR coupling strengths $g_{r}(\xi)$ and $g_{cr}(\xi)$ via the dimensionless modulation amplitude $\xi$. To make Hamiltonian (12) time independent, we introduce the effective frequency for the qubit
$\tilde{\omega}_{0}$ $=\frac{\Delta_{m_{0}}+\delta_{n_{0}}}{2},$
and the cavity field
$\tilde{\omega}_{c}=\frac{\Delta_{m_{0}}-\delta_{n_{0}}}{2}.$
After transferring to the second rotating frame by the unitary transformation
@@
Here, we have introduced the anisotropic parameter $\epsilon(\xi)=g_{cr}(\xi)/g_{r}(\xi)$. The effective coupling strength $g_{r}(\xi)$, the effective frequencies $\tilde{\omega}_{0}$ and $\tilde{\omega}_{c}$, and anisotropic parameter $\epsilon(\xi)$ can be regulated by the modulation parameters $\nu$ and $\xi$. The effective Hamiltonian (15) describe an anisotropic QRM with $\epsilon(\xi)\neq 1$ and a standard QRM with $\epsilon(\xi)=1$. Hereafter, we write $\epsilon(\xi)$ as $\epsilon$ for short.
@@ III Ultrastrong coupling
The effective standard quantum Rabi model ($\epsilon=1$) in the limit $\tilde{\eta}=\tilde{\omega}_{0}/\tilde{\omega}_{c}\rightarrow\infty$ can exhibit quantum phase transition from the normal phase to the superradiant phase. This phase transition occurs at $g_{C}=\sqrt{\tilde{\omega}_{0}\tilde{\omega}_{c}}/2$ Ashhab2013 above which $Z_{2}$ parity symmetry is broken and the bosonic mode is macroscopically occupied Hwang2015Rabi . Here, the critical coupling strength $g_{C}$ depends on the modulation frequency $\nu$. Since the occurrence of the QPT in the quantum Rabi model requires the limit $\tilde{\eta}\rightarrow\infty$, to see clearly the dependence of $\tilde{\eta}$ on the modulation parameters, we plot $\tilde{\eta}$ versus $\nu$ for different ratios $\eta=\omega_{0}/\omega_{c}^{\prime}$ in Fig. 2(a). $\tilde{\eta}$ is independent of $\xi$. As shown in Fig. 2(a), the curve of the effective ratio $\tilde{\eta}$ is composed by series of V-type valleys until $\nu$ increases up to $\nu/\omega_{0}=2(\omega_{0}+\omega_{c}^{\prime})$. The V-shaped valleys share the same $m_{0}$ and the valley for $m_{0}$ starts at $\nu=2(\omega_{0}+\omega_{c}^{\prime})/(2\left|m_{0}\right|+1)$. When $\nu>2(\omega_{0}+\omega_{c}^{\prime})$, $m_{0}=0$ and $\tilde{\eta}=\eta$. Thus, in the regime $\nu<2(\omega_{0}+\omega_{c}^{\prime})$, we can manipulate $\tilde{\eta}$ via choosing appropriate $\nu$ while in the regime $\nu>2(\omega_{0}+\omega_{c}^{\prime})$, this $\tilde{\eta}$ can not be varied via changing $\nu$. To see clearly the dependence of the effective coupling strengths $g_{r}(\xi)$ and $g_{cr}(\xi)$ on the modulation amplitude $\xi$, we plot $g_{cr}/g_{C}$ and $g_{r}/g_{C}$ versus the modulation amplitude $\xi$ for $\eta=1$ and $10^{2}$ at $\nu/\omega_{0}=0.49$ and $0.68$ in Figs. 2(b) and 2(c), respectively. The effective coupling strength $g_{r}/g_{C}\ (g_{cr}/g_{C})$ depends on the modulation amplitude $\xi$ in the form of the Bessel function. When $\nu/\omega_{0}=0.49$ for $\eta=1$ $(m_{0}=-4,\ n_{0}=-1)$, $g_{r}/g_{C}$ oscillates between -2.41656 and 6 with reducing amplitude, and $g_{cr}/g_{C}$ oscillates between -1.64629 and 2.39791, as shown in Fig. 2(b). When $\nu/\omega_{0}=0.68$ for $\eta=10^{2}$ $(m_{0}=n_{0}=-1)$, $g_{r}/g_{C}$ and $g_{cr}/g_{C}$ show the same oscillating feature between -1.23432 and 0.734245, as shown in Fig. 2(c). Note that, $g_{r}=g_{cr}$ in Fig. 2(c) under the chosen parameters. The results show that one can tune the effective coupling strengths $g_{r}$ and $g_{cr}$ in a broad range.
Figure 2: (Color online) (a) The effective ratio $\tilde{\eta}$ versus the modulation frequency $\nu/\omega_{0}$ for $\eta=10^{2}\text{ and }10^{3}.$ (b-c) The couplings $g_{cr}/g_{C}$ and $g_{r}/g_{C}$ versus the modulation amplitude $\xi$. Here, $\nu/\omega_{0}=0.49$ and $\eta=1$ in panel (b) and $\nu/\omega_{0}=0.68$ and $\eta=10^{2}$ in panel (c), other parameter is $g/\omega_{0}=0.06.$

### v2

# Controlling the superradiant phase transition in the quantum Rabi model

@@ Abstract
In the ultrastrong-coupling regime, the quantum Rabi model can exhibit quantum phase transition (QPT) when the ratio of the qubit transition frequency to the frequency of the cavity field approaches infinity. However, it is challenging to control the QPT in few-body systems because of the limited coupling strength and the $A^{2}$ term. Here, we propose a practical scheme to manipulate the QPT of quantum Rabi model in the strong-coupling regime. By applying a periodic frequency modulation to the two-level system in a quantum Rabi model in the strong-coupling regime, an anisotropic quantum Rabi model with ultrastrong and tunable coupling strengths for rotating and counter-rotating terms is obtained. The ground state and excitation energy of this model, in terms of the modulation parameters, are studied. We find that the QPT of the quantum Rabi model can be observed in the strong-coupling regime and externally controlled by the modulation.
@@ I Introduction
Quantum phase transition (QPT) QPT1999 has long been an important and fundamental phenomenon in many-body systems, and has attracted a remarkable amount of study interest due to its special features. The QPT occurs when the non-thermal parameter, such as coupling strength, exceeds a critical value at zero temperature QPT1999 . For example, the superradiant phase transition occurs in the Dicke model Dicke when the light-matter coupling strength exceeds a critical value in the thermodynamic limit, i.e., $N\rightarrow\infty$ 1973dickle ; 1973spt ; Sunchangpu ; liyong ; CEPRL ; NLPRL ; ABPRL ; PRL120183603 . The control of the QPT in Dicke model was discussed ABPRL ; 2023Modulation . Recently, it was found that the QPT can also occur in few-body systems, such as the quantum Rabi model (QRM) 1936Rabi ; 1937Rabi when the ratio of the qubit transition frequency $\omega_{0}$ to the frequency $\omega_{c}$ of the cavity field tends to infinity, i.e., $\eta=\omega_{0}/\omega_{c}\rightarrow\infty$ Hwang2015Rabi . Hereafter, many efforts were devoted to studying the QPT in QRM both theoretically MaoXinLiu2017 ; Rabi triangle ; Rabi lattice ; dispassive QRM ; dar ; sar ; yingzujian ; PRAaqrm ; L. T. ShenPRA ; full quantum limit and experimentally trapped ions ; nuclear ; ZhengShiBiao , such as the universal scaling and critical exponents MaoXinLiu2017 , QPT in the quantum Rabi triangle Rabi triangle ; Rabi lattice , dissipative phase transition in the open QRM dispassive QRM ; dar ; sar , QPT in the anisotropic QRM yingzujian ; PRAaqrm and quench dynamics in the anisotropic QRM L. T. ShenPRA . Even the QPT in QRM in the fully quantum limit has also been predicted full quantum limit .
@@
Though the QPT in QRM was experimentally observed in a single trapped ion trapped ions and a nuclear magnetic resonance quantum simulator nuclear , it is still challenging to manipulate the ground-state QPT due to the limited coupling strength with the classical field (dispersion) limit $\eta\rightarrow\infty$ and the $A^{2}$ term in cavity quantum electrodynamics (QED) A2 term ; Hwang2015Rabi ; no-go theorm . The $A^{2}$ term can prevent the occurrence of the QPT in QRM Hwang2015Rabi , which can not be neglected in the ultrastrong-coupling regime in cavity-QED system.
@@
In this work, we propose a practical scheme to manipulate the QPT in QRM. Concretely, a periodic modulation is applied to the qubit transition frequency in a standard QRM 2017manipulation ; C. Liu in the strong-coupling regime. Then we obtain an anisotropic QRM with tunable rotating and counter-rotating coupling strengths. Through the modulation, both the effective rotating and counter-rotating (CR) coupling strengths can be varied in a wide range and reach the ultrastrong-coupling regime. The ground state is allowed to be in the superradiant phase. To investigate the QPT of the anisotropic QRM, we calculate excitation energy and the ground-state phase in both the normal and superradiant phase in the entire coupling parameter space. We find that both the normal and superradiant phase can be reached in the strong-coupling regime via the modulation. We also find that Goldstone modes Hwang2016JC ; Goldstone exist in the superradiant phase. The dependence of excitation energy and ground-state phase on the modulation parameters is analyzed. The feature of the QPT is also investigated by calculating the derivatives of the ground-state energy. Our results show that the QPT in the QRM can be observed and manipulated externally in the strong-coupling regime. These results may inspire other studies on QPT engineering in other few-body systems.
@@
The rest of this paper is organized as follows. In Sec. II, we introduce a qubit frequency modulation in a standard QRM and obtain an effective anisotropic QRM whose rotating and counter-rotating terms are tunable. In Sec. III, we analyze the dependence of the effective frequency ratio and the effective coupling strengths on modulation parameters. We obtain the excitation energies for five phases and the phase boundary in Sec. IV. In Sec. V, we figure out the phase diagram and verify it by analyzing the order parameters. The feature of QPT is also discussed. We also discuss how to manipulate the ground-state phase by varying the modulation parameters in Sec. VI. In Sec. VII, we discuss the $A^{2}$ term under modulation. We present some discussions concerning the experimental implementation in Sec. VIII. A brief conclusion is given in Sec. IX. Finally, three Appendixes are presented to show the validity evaluation of the rotating-wave approximation, the operators for superradiant phase, and the squeezing parameters.
@@ II Model and Hamiltonian
The system we consider is composed of a two-level system (qubit) coupled to a cavity field, as shown in Fig. 1, which is described by the quantum Rabi model 1936Rabi ; 1937Rabi . The corresponding Hamiltonian reads
@@
is the Jaynes-Cummings (JC) Hamiltonian, and
@@
describes the CR interactions. Here, $\sigma_{\pm}=(\sigma_{x}\pm i\sigma_{y})$/2 with $\sigma_{x,y,z}$ being the Pauli matrices, and $a\ (a^{\dagger})$ is the annihilation (creation) operator of the cavity field with frequency $\omega_{c}$. We use $\left|e\right\rangle$ and $\left|g\right\rangle$ to label the excited and ground states of the qubit, respectively. The transition frequency of the qubit is $\omega_{0}$ and the coupling strength between the qubit and the cavity field is $g$. Note that in this work we use the terms of the cavity-QED system to expound the physical idea. Actually, the physical results also work for other physical platforms, in which the physical model consisting of a two-level system and a single-mode bosonic field can be implemented.
@@
where $\xi$ and $\nu$ are the dimensionless modulation amplitude and frequency, respectively. After including the two-photon $A^{2}$ term
@@
$V_{1}(t)$ $=\exp\left\{-i\left[\left(\omega_{c}^{\prime}a^{\dagger}a+\frac{1}{2}\omega_{0}\sigma_{z}\right)t+\frac{1}{2}\xi\sin(\nu t)\sigma_{z}\right]\right\},$ (8)
the Hamiltonian becomes
@@
where we omit the constant term in $H_{A^{2}}$. Here, $\delta_{n}=\omega_{0}-\omega_{c}^{\prime}+n\nu$ is the oscillating frequency for the $n$th sideband in the rotating terms, and $\Delta_{m}=\omega_{0}+\omega_{c}^{\prime}+m\nu$ is the oscillating frequency for the $m$th sideband in the CR terms.
In the derivation of $\tilde{H}(t)$, we use the Jacobi-Anger expansion
$e^{i\xi\sin(\nu t)}=\sum_{n=-\infty}^{\infty}J_{n}(\xi)e^{in\nu t},$ (10)
where $J_{n}(\xi)$ is the $n$th Bessel function of the first kind. In Eq. (9), a rotating sideband $n_{0}$ and a CR sideband $m_{0}$ satisfying $\left|gJ_{n_{0}}(\xi)/\delta_{n_{0}}\right|=\max\{\left|gJ_{n}(\xi)/\delta_{n}\right|,n\in Z\}$ and $\left|gJ_{m_{0}}(\xi)/\Delta_{m_{0}}\right|=\max\{\left|gJ_{m}(\xi)/\Delta_{m}\right|,m\in Z\}$ can be selected. Here, “Z” denotes all integers. The oscillating frequencies of all rotating and CR sidebands can be written as $\delta_{n_{0}+s}=\delta_{n_{0}}+s\nu$ and $\Delta_{m_{0}+s}=\Delta_{m_{0}}+s\nu$ with $s$ an integer, respectively. Thus the other sidebands are separated from the $n_{0}$ ($m_{0}$) sideband by $s\nu$. Under the conditions
@@
all the sidebands other than the $n_{0}$($m_{0}$)th sideband can be discarded using the rotating-wave approximation (RWA). With $g_{A^{2}}\ll g\ll\omega_{0}$ and $g_{A^{2}}\ll 2\omega_{c}^{\prime}$, the $a^{2}$ and $a^{\dagger 2}$ terms with oscillating frequencies ±$\pm 2\omega_{c}^{\prime}$ from the two-photon $A^{2}$ Hamiltonian can also be neglected using the RWA. To be precise, the validity of this approximation is discussed in detail in Sec. VII.
The Hamiltonian $\tilde{H}(t)$ then can be approximated as
@@
where we introduce the effective coupling strengths
$g_{r}(\xi)=gJ_{n_{0}}(\xi),\hskip 14.22636ptg_{cr}(\xi)=gJ_{m_{0}}(\xi).$ (13)
To evaluate the validity of the RWA made in Hamiltonian (9), we examine the dynamics of the fidelity between the evolving states governed by Hamiltonians (9) and (12). The results are shown in Appendix A. The Hamiltonian describes an effective standard QRM with $g_{r}=g_{cr}$ or an effective anisotropic QRM with $g_{r}\neq g_{cr}$ QTxie in interaction picture. Note that, we can tune the effective rotating and CR coupling strengths $g_{r}(\xi)$ and $g_{cr}(\xi)$ via the dimensionless modulation amplitude $\xi$.
To make Hamiltonian (12) time independent, we introduce the effective frequencies
[…]
for the qubit and the cavity field, respectively. Here, $\tilde{\omega}_{0}(\nu)$ and $\tilde{\omega}_{c}(\nu)$ depend on the modulation frequency $\nu$ via $\Delta_{m_{0}}$ and $\delta_{n_{0}}$. Hereafter, we denote $\tilde{\omega}_{0,c}(\nu)$ as $\tilde{\omega}_{0,c}$ for precision. After transferring to the second rotating frame by the unitary transformation
@@
Here, we introduce the anisotropic parameter $\epsilon(\xi)=g_{cr}(\xi)/g_{r}(\xi)$. The effective coupling strength $g_{r}(\xi)$, the effective frequencies $\tilde{\omega}_{0}(\nu)$ and $\tilde{\omega}_{c}(\nu)$, and the anisotropic parameter $\epsilon(\xi)$ can be tuned by the modulation parameters $\nu$ and $\xi$. The effective Hamiltonian (16) describe an anisotropic QRM with $\epsilon(\xi)\neq 1$ and a standard QRM with $\epsilon(\xi)=1$. Hereafter, we denote $\epsilon(\xi)$ as $\epsilon$ for precision.
Figure 2: (a) The effective ratio $\tilde{\eta}$ versus the modulation frequency $\nu/\omega_{0}$ for $\eta=10^{2}\text{ and }10^{3}.$ (b), (c) The couplings $g_{cr}/g_{C}$ and $g_{r}/g_{C}$ versus the modulation amplitude $\xi$. Here, $\nu/\omega_{0}=0.49$ and $\eta=1$ in panel (b) and $\nu/\omega_{0}=0.68$ and $\eta=10^{2}$ in panel (c), the other parameter is $g/\omega_{0}=0.06.$
@@ III Ultrastrong coupling
The effective QRM ($\epsilon=1$) in the limit $\tilde{\eta}=\tilde{\omega}_{0}/\tilde{\omega}_{c}\rightarrow\infty$ can exhibit quantum phase transition from the normal phase to the superradiant phase. This phase transition occurs at $g_{C}=\sqrt{\tilde{\omega}_{0}\tilde{\omega}_{c}}/2$ Ashhab2013 above which $Z_{2}$ parity symmetry is broken and the bosonic mode is macroscopically occupied Hwang2015Rabi . Here, the critical coupling strength $g_{C}$ depends on the modulation frequency $\nu$. Since the occurrence of the QPT in the quantum Rabi model requires the limit $\tilde{\eta}\rightarrow\infty$, to see clearly the dependence of $\tilde{\eta}$ on the modulation parameters, we plot $\tilde{\eta}$ versus $\nu$ for different ratios $\eta=\omega_{0}/\omega_{c}^{\prime}$ in Fig. 2(a). $\tilde{\eta}$ is independent of $\xi$. As shown in Fig. 2(a), the curve of the effective ratio $\tilde{\eta}$ is composed of a series of V-type valleys until $\nu$ increases up to $\nu/\omega_{0}=2(\omega_{0}+\omega_{c}^{\prime})$. The V-shaped valleys share the same $m_{0}$ and the valley for $m_{0}$ starts at $\nu=2(\omega_{0}+\omega_{c}^{\prime})/(2\left|m_{0}\right|+1)$. When $\nu>2(\omega_{0}+\omega_{c}^{\prime})$, $m_{0}=0$ and $\tilde{\eta}=\eta$. Thus, in the regime $\nu<2(\omega_{0}+\omega_{c}^{\prime})$, we can manipulate $\tilde{\eta}$ via choosing appropriate $\nu$ while in the regime $\nu>2(\omega_{0}+\omega_{c}^{\prime})$, this $\tilde{\eta}$ cannot be varied via changing $\nu$. As shown in Fig. 2(a), to satisfy the condition (11) and the limit $\tilde{\eta}\rightarrow\infty$, the value of $\nu$ can be chosen in a broad range.
To see clearly the dependence of the effective coupling strengths $g_{r}(\xi)$ and $g_{cr}(\xi)$ on the modulation amplitude $\xi$, we plot $g_{cr}/g_{C}$ and $g_{r}/g_{C}$ versus the modulation amplitude $\xi$ for $\eta=1$ and $10^{2}$ at $\nu/\omega_{0}=0.49$ and $0.68$ in Figs. 2(b) and 2(c), respectively. The effective coupling strength $g_{r}/g_{C}\ (g_{cr}/g_{C})$ depends on the modulation amplitude $\xi$ in the form of the Bessel function. When $\nu/\omega_{0}=0.49$ for $\eta=1$ $(m_{0}=-4,\ n_{0}=-1)$, $g_{r}/g_{C}$ oscillates between $-2.41656$ and 6 with reducing amplitude, and $g_{cr}/g_{C}$ oscillates between $-1.646$ and 2.398, as shown in Fig. 2(b). When $\nu/\omega_{0}=0.68$ for $\eta=10^{2}$ $(m_{0}=n_{0}=-1)$, $g_{r}/g_{C}$ and $g_{cr}/g_{C}$ show the same oscillating feature between $-1.234$ and 0.734, as shown in Fig. 2(c). The results show that one can tune the effective coupling strengths $g_{r}$ and $g_{cr}$ in a broad range.
Figure 3: The phase transition boundary $\tilde{g}_{C}/\omega_{0}$ versus $\xi$ for the anisotropic QRM at $\eta=10^{2}$ and $\nu/\omega_{0}=0.402$, $0.68$, and $0.8$.